Find the area of an equilateral triangle that has a perimeter of 21 inches. Round the answer to one decimal place. Thank you.
Answer by jessica43(140)

An equilateral triangle is a triangle with all three sides of equal length. So if we know the perimeter of a triangle is 21 inches, we can divide that by 3 to find the length of each side:
21/3 = 7
So each side of the triangle is 7 inches.
Now we will use the area of an equilateral triangle formula to solve the problem:
A = ((s^2)*(sqrt(3)))/4 where s = the length of a side
A = ((7^2)*(sqrt(3)))/4
A = (49*(sqrt(3)))/4
A = 21.2
So the area of the equilateral triangle is 21.2 square inches.